OpenStudy (anonymous):

3 more than 5 times a number is equal to 7 more than the number. Find the number.

OpenStudy (anonymous):

more means + is = number x

OpenStudy (anonymous):

i wrote it as 3+5*x=7+x

OpenStudy (anonymous):

perfect, now solve the equation to find x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what's the number?

OpenStudy (anonymous):

idk i'm confused on how to begin solving it,

OpenStudy (anonymous):

first take the x's to one side of the equation and the numbers to the other,don't forget to flip signs

OpenStudy (anonymous):

ok i understand bringing the x's to one side but there's a times x & a +x how do i work with that?

OpenStudy (anonymous):

3+5x=7+x 5x-x=7-3 got it?

OpenStudy (anonymous):

ohhh i wrote it wrong! i had 3+5*x=7+x

OpenStudy (anonymous):

5*x is 5x

OpenStudy (anonymous):

ok i got x=1

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

THX 4 the help!!

OpenStudy (anonymous):

yw